The management of Ro Corporation is investigating automating a process. Old equipment, with a current salvage value of $12,000, would be replaced by a new machine. The new machine would be purchased for $444,000 and would have a 6 year useful life and no salvage value. By automating the process, the company would save $153,000 per year in cash operating costs. The simple rate of return on the investment is closest to (Ignore income taxes.):

Multiple Choice

  • 18.3%

  • 17.8%

  • 34.5%

  • 16.7%
Answer #1

The correct answer is 17.8%

Calculation

Net impact on net income = saving in cash operating costs- depreciation

= 153000 - (444000/6)

= 153000 - 74000

= $ 79000

Simple rate of return = net impact on net income ÷ investment

=79000÷444000

=17.8%

Thus the correct answer is 17.8%